OpenStudy (anonymous):

Find the missing digit to make 18,59_ divisible by 9

OpenStudy (cwrw238):

18 / 9 = 2 59 / 9 = 6 remainder 5 5x must be divisible by 9 what is x?

OpenStudy (btaylor):

the sum of all the digits must be a multiple of 9. 1+8+5+9 = 23, so the missing digit must be 4 in order to make it divisible by 9

OpenStudy (cwrw238):

ah

OpenStudy (anonymous):

thanks, so add them all, then find a number that leaves it a whole number, not a decimal?

OpenStudy (btaylor):

, no, find a number that makes it a multiple of 9 (9, 18, 27, 36, 45...)

OpenStudy (cwrw238):

add a number so its a multiple of 9 23 + 4 = 27 which is a multiple of 9

OpenStudy (anonymous):

OOohh, okay. :D